OpenStudy (anonymous):

which expression is equivalent to (-6p+7)(-4)

OpenStudy (vishweshshrimali5):

Just multiply to get req. expression

OpenStudy (lgbasallote):

\[(-6p+7)(-4) \implies (-6p \times -4) + (7\times -4)\] does that help?

OpenStudy (anonymous):

\[(-a + b) (-c) = ac - bc\]

OpenStudy (lgbasallote):

@footballkidd12 do you get these?